OpenStudy (anonymous):

Write the point-slope form of the equation passing through (5, -1) with a slope of 6.

OpenStudy (solomonzelman):

point slope, where slope is m, and point is \(\large\color{black}{(x_1,y_1) }\) \(\large\color{black}{ y-y_1=m(x-x_1) }\)

OpenStudy (solomonzelman):

Examples: ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~ when slope is 3, and point is \(\large\color{black}{ (4,5) }\) then, the line is: \(\large\color{black}{ y-5=3(x-4) }\) ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~ when slope is 2, and point is \(\large\color{black}{ (-8,1) }\) then, the line is: \(\large\color{black}{ y-1=2(x+8) }\) ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~ when slope is 5, and point is \(\large\color{black}{ (2,-3) }\) then, the line is: \(\large\color{black}{ y+3=5(x-2) }\)

OpenStudy (nincompoop):

I love what you are doing, @SolomonZelman Spread the ways of teaching proper concept

OpenStudy (solomonzelman):

tnx, lol:) trying my best...

OpenStudy (solomonzelman):

but you know if I keep disconnecting ever other question it is very discouraging,,,,

OpenStudy (solomonzelman):

anyway, do you have anything to tell me readhed, tnx, concerns, or something?

OpenStudy (anonymous):

ok thank you i for your help:)

OpenStudy (solomonzelman):

Sure. (Just got my connection back) can you tell me the answer you got, so that I can see it?